# Transient and steady-state chaos in dissipative quantum systems

### v1

@@ Abstract
Dissipative quantum chaos plays a central role in the characterization and control of information scrambling, non-unitary evolution, and thermalization, but it still lacks a precise definition. The Grobe-Haake-Sommers conjecture, which links Ginibre level repulsion to classical chaotic dynamics, was recently shown to fail [Phys. Rev. Lett. 133, 240404 (2024)]. We properly restore the quantum-classical correspondence through a dynamical approach based on entanglement entropy and out-of-time-order correlators (OTOCs), which reveal signatures of chaos beyond spectral statistics. Focusing on the open anisotropic Dicke model, we identify two distinct regimes: transient chaos, marked by rapid early-time growth of entanglement and OTOCs followed by low saturation values, and steady-state chaos, characterized by high long-time values. We introduce a random matrix toy model and show that Ginibre spectral statistics signals short-time chaos rather than steady-state chaos. Our results establish entanglement dynamics and OTOCs as reliable diagnostics of dissipative quantum chaos across different timescales.
@@
Quantum chaos is connected to a wide range of theoretical and experimental phenomena, from thermalization [1, 2] and the failure of many-body localization [3], to the scrambling of quantum information [4, 5] and the exponential growth of perturbations near black hole horizons [6]. While chaos in closed quantum systems is typically characterized by random matrix theory [7, 8], including Wigner-Dyson level statistics [9, 10, 11] and eigenstates resembling random vectors [12, 1], defining chaos in open quantum systems remains a challenge [13, 14]. Dissipation alters dynamical behavior, raising questions about the quantum-classical correspondence in non-unitary settings.
@@
Efforts to define quantum chaos in open systems have led to the Grobe-Haake-Sommer conjecture [15, 16, 10, 17, 18, 19, 20, 21, 22, 23, 24, 25], which relates classical chaotic attractors with Ginibre level repulsion in the spectrum of the Liouvillian superoperator. However, recent studies have shown that this correspondence can fail [26, 27, 28]. In particular, Ginibre spectral statistics may emerge even when the long-time dynamics is regular [26]. This discrepancy suggests that spectral indicators alone may be insufficient to capture the full nature of dissipative quantum chaos.
In this work, we adopt a dynamical perspective and demonstrate that the quantum-classical correspondence can be restored when chaos is diagnosed through the evolution of entanglement and out-of-time-order correlators (OTOCs). Unlike closed Hamiltonian systems, where rapid early-time mixing often indicates persistent chaotic dynamics, dissipation can suppress long-time chaos, despite signs of chaotic behavior at short times. As illustrated in Fig. 1, we identify three regimes with qualitatively distinct behavior: (I) steady-state chaos, marked by fast initial entropy growth and large long-time saturation values of entanglement and OTOCs (middle panel), in agreement with the presence of a chaotic attractor in the classical dynamics (see Bloch sphere on the left side); (II) transient chaos, which also features rapid early-time scrambling but evolves into a low-entanglement steady state, consistent with the presence of a stable attractor; and (III) regular regime, with slow dynamics and no chaos at any time. Importantly, the right panels of Fig. 1 show that both transient and steady-state chaos exhibit Ginibre level statistics in the Liouvillian spectrum, demonstrating the limitations of spectral features in fully characterizing dissipative quantum chaos.
@@
Figure 1: Schematic illustration. A dynamical framework restores the quantum-classical correspondence by identifying three regimes: (I) steady-state chaos, featuring a chaotic attractor in the classical dynamics (left), and linear growth followed by large saturation of entanglement entropy (middle panel); (II) transient chaos, characterized by a regular attractor, and rapid short-time entanglement growth followed by decay at long times; and (III) regular dynamics, showing slow entanglement growth and low saturation. Regimes (I) and (II) exhibit Ginibre Liouvillian spectral statistics (right panels), while (III) corresponds to 2d-Poisson statistics.
@@
We explore these ideas using the anisotropic Dicke model (ADM) [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50] in the presence of photon loss. We show that, unlike Liouvillian spectral statistics [26], the dynamics of entanglement entropy (EE) and OTOCs distinguish transient from steady-state chaos, in agreement with classical diagnostics based on Lyapunov exponents.
@@
To support and generalize our findings, we introduce a random matrix toy model with a tunable Liouvillian. We show that Ginibre spectral statistics emerge whenever the entanglement entropy exhibits rapid initial linear growth, regardless of whether it saturates at high values (steady-state chaos) or decays to low values (transient chaos). This analysis reveals that Liouvillian spectral statistics can reflect early-time chaotic behavior, but does not distinguish different asymptotic regimes.
@@
^$\hat{\mathcal{H}}\!=\omega\hat{a}^{\dagger}a\!+\!\omega_{0}\hat{S}_{z}\!+\!\frac{\lambda_{-}}{\sqrt{2S}}(\hat{a}\,\hat{S}_{+}\!+\!\hat{a}^{\dagger}\hat{S}_{-})\!+\!\frac{\lambda_{+}}{\sqrt{2S}}(\hat{a}\hat{S}_{-}\!+\!\hat{a}^{\dagger}\hat{S}_{+}),$
where $\hbar=1$, $\hat{a}$ ($\hat{a}^{\dagger}$) annihilates (creates) a cavity photon mode with frequency $\omega$, the collective pseudospin operators $\hat{S}_{z,+,-}$ describe the joint behavior of $N$ two-level atoms with energy splitting $\omega_{0}$, and $\lambda_{\pm}$ are the atom-photon coupling strengths. In all figures, we set $S=N/2=5$, $\omega=1.0$, and $\omega_{0}=1.0$.
@@
$\frac{d\hat{\rho}}{dt}=\hat{{\cal L}}[\hat{\rho}]=-i[\hat{\mathcal{H}},\hat{\rho}]+\kappa\left[2\hat{a}\hat{\rho}\hat{a}^{\dagger}-\left\{\hat{a}^{\dagger}\hat{a},\hat{\rho}\right\}\right],$ (1)
@@
The classical dynamics is derived from the quantum evolution equation, $d\langle\hat{O}\rangle/dt={\rm Tr}(\hat{O}\dot{\hat{\rho}})$, in the limit of a large collective spin, $S\rightarrow\infty$. A phase diagram of the asymptotic classical dynamics of the open ADM was obtained in [37] by analyzing the stability of the fixed points of the equations of motion. The system exhibits a variety of dynamical phases, including normal, superradiant, limit-cycle, and chaotic regimes [see Supplemental Material (SM) [54]].
@@
Steady-state chaos.– In the open classical ADM, chaotic dynamics emerges when the regular attractors become unstable. In the quantum domain, the steady state is defined by the zero eigenvalue of the Liouvillian superoperator. We show that this long-time quantum state can retain signatures of the underlying classical chaos, a phenomenon we refer to as steady-state chaos.
@@
Since in isolated quantum systems, chaos is typically associated with enhanced entanglement [2, 1, 55, 56, 57, 58], we investigate the time evolution of the EE in the presence of dissipation. The system is initially prepared in a product state consisting of a spin coherent state and a photon coherent state [59]. Entanglement builds up dynamically due to the atom-photon interaction. We analyze the total entanglement entropy, defined as $\mathcal{S}^{\rm en}(t)=\mathcal{S}_{\rm spin}^{\rm en}(t)+\mathcal{S}_{\rm photon}^{\rm en}(t)$, where $\mathcal{S}_{\rm spin}^{\rm en}$ ($\mathcal{S}_{\rm photon}^{\rm en}$) is the entropy of the reduced spin (photon) subsystem obtained by tracing out the photon (spin) degrees of freedom.
@@
Contrary to Hamiltonian systems, where the long-time saturation value of EE depends on the initial state, dissipation drives the system toward a unique asymptotic behavior. In dissipative dynamics, only the short-time evolution of the EE depends on the initial state.
@@
Figure 2: Signature of steady-state chaos on the $\lambda_{-}-\lambda_{+}$ plane. (a), (c) Averaged long-time Lyapunov exponent $\Lambda_{\rm ss}$ and (b), (d) total entanglement entropy $\mathcal{S}_{\rm ss}^{\rm en}$ in the (a)-(b) absence of dissipation, $\kappa=0$, and (c)-(d) presence of dissipation, $\kappa=1$.
@@
The color plots in Fig. 2 compare the average long-time Lyapunov exponent [60, 61, 62], denoted by $\Lambda_{\rm ss}$ [Figs. 2(a),2(c)], and the steady-state EE, $\mathcal{S}_{\rm ss}^{\rm en}$, averaged over an ensemble of initial states [Figs. 2(b),2(d)] for the isolated [Figs. 2(a)-(b)] and dissipative [Figs. 2(c)-(d)] cases. In the absence of dissipation, the system exhibits large values of $\Lambda_{\rm ss}$ and $\mathcal{S}_{\rm ss}^{\rm en}$ over a broad region of the coupling parameter space, especially near the Dicke limit, $\lambda_{-}=\lambda_{+}=\lambda$, where classical dynamics is maximally chaotic. When dissipation is introduced, the extent of the chaotic region is significantly reduced. As seen in Figs. 2(c)-(d), large values of $\Lambda_{\rm ss}$ and $\mathcal{S}_{\rm ss}^{\rm en}$ persist only within a narrow triangular region of the parameter space. Outside this region, the classical system becomes regular [see Fig. 2(c)], including the area around the Dicke limit, and in the quantum case, small nonzero values of $\mathcal{S}_{\rm ss}^{\rm en}$ remain visible [Fig. 2(d)], due to the influence of quantum fluctuations at finite spin magnitude $S$ and the proximity to a non-equilibrium phase transition.
@@
The key result in Fig. 2 is the strong quantum-classical correspondence observed not only in the isolated regime but also in the presence of dissipation. The saturation value of the EE closely follows the classical Lyapunov exponent, as shown by the close agreement between Figs. 2(c) and 2(d). This demonstrates that, unlike Liouvillian level statistics, which can present Ginibre spectral correlations when the open classical model is regular [26], the steady-state EE provides a reliable quantum signature of dissipative classical chaos.
@@
In isolated systems, the OTOC has also been extensively studied as a quantum diagnostic of chaos [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 39, 40, 76, 77]. It is defined as $\mathcal{F}(t)=\langle\hat{W}^{\dagger}(t)\hat{V}^{\dagger}(0)\hat{W}(t)\hat{V}(0)\rangle$, where the operator $\hat{W}$ evolves in time and the operator $\hat{V}$ is fixed at $t=0$. The OTOC quantifies the spread of quantum information and sensitivity to perturbations. A variant of this correlator is the fidelity OTOC (FOTOC) [75] obtained by choosing $\hat{V}=\hat{\rho}(0)$ and $\hat{W}=e^{i\delta\phi\hat{G}}$, where $\delta\phi$ is a small perturbation and $\hat{G}$ is a Hermitian operator. In the perturbative limit $\delta\phi\ll 1$, the FOTOC can be written in terms of the variance $(\Delta{G}(t))^{2}=\langle\hat{G}^{2}(t)\rangle-\langle\hat{G}(t)\rangle^{2}$ as $\mathcal{F}_{G}(t)\approx 1-\delta\phi^{2}(\Delta{G}(t))^{2}$.
@@
We generalize the FOTOC to open quantum systems by evolving the operator $\hat{W}$ under non-unitary dynamics according to $d\hat{W}/dt=\mathcal{L}^{\dagger}[\hat{W}]$. Similar to the EE, the steady-state value of the standard deviation $(\Delta G)_{\rm ss}$ obtained from the FOTOC also serves as an indicator of steady-state chaos (see SM [54]).
@@
Transient chaos.– Complementing the long-time analysis, we examine the short-time dynamics of the EE and $\Delta G(t)$, both of which capture signatures of transient chaos. In isolated classical systems, chaos arises from rapid mixing and sensitivity to initial conditions that emerge early in the evolution. When dissipation is introduced, the system may still display chaotic behavior at short times, even if the long-time dynamics become regular. This transient chaos reflects a competition between the early-time chaotic dynamics inherited from the isolated system and the eventual convergence to a regular attractor of the dissipative system. In the quantum regime, a similar phenomenon occurs: the dynamics can ultimately settle into a non-chaotic steady state, even if remnants of chaos, such as rapid entropy growth, are still visible at early times. In our system, the Dicke limit offers a natural setting to explore this behavior, as it exhibits maximal chaos in the absence of dissipation, while increasing the dissipation rate $\kappa$ progressively suppresses chaos at long times, as the superradiant fixed point becomes a stable attractor.
@@
Figure 3: Transient chaos vs steady-state chaos in the Dicke limit $\lambda_{-}=\lambda_{+}=\lambda$. (a) Finite-time ensemble averaged Lyapunov exponent $\Lambda_{t}$, (b) total entanglement entropy $\mathcal{S}^{\rm en}(t)$, and (c) standard deviation $\Delta S_{z}(t)$ for atom-photon coupling strengths $\lambda=0.8,1.5$ and $\kappa=0,1$. (d) The growth rate of $\mathcal{S}^{\rm en}(t)$ and the time-averaged Lyapunov exponent $\overline{\Lambda}$ within a time interval $t\in[0,0.5]$ as a function of the coupling strength $\lambda$ for $\kappa=1$. The inset of (d) shows the initial linear growth of $\mathcal{S}^{\rm en}(t)$ for $\lambda=1.5$ and $\kappa=1$.
@@
To uncover quantum signatures of transient chaos in the Dicke limit, we examine the time evolution of the entanglement entropy $\mathcal{S}^{\rm en}(t)$ [Fig. 3(b)] and the standard deviation of the $z$ component of the spin operator, $\Delta S_{z}(t)$ [Fig. 3(c)], both averaged over an ensemble of initial states. In both isolated and dissipative settings, these observables exhibit rapid growth at short times, consistent with short-time chaotic dynamics. However, at long times, similar to the behavior of $\Lambda_{t}$, the asymptotic values of $\mathcal{S}^{\rm en}(t)$ and $\Delta S_{z}(t)$ are significantly reduced for the dissipative model, indicating a transition to regular dynamics. This contrast between rapid short-time growth and long-time suppression reflects the presence of transient chaos also for the quantum dissipative system. Similar behavior is also observed for the standard deviation of the photon number.
Further insight into the nature of transient chaos is provided by its dependence on the atom-photon coupling $\lambda$. As seen in Figs. 3(b)-(c), the short-time growth rates of $\mathcal{S}^{\rm en}(t)$ and $\Delta S_{z}(t)$ increase with the atom-photon coupling $\lambda$, in agreement with the larger values of the short-time Lyapunov exponent observed in Fig. 3(a). This dependence on $\lambda$ is quantified in Fig. 3(d), where we show that both the growth rate of $\mathcal{S}^{\rm en}(t)$ and the time-averaged Lyapunov exponent computed over small time windows $t\in[0,0.5]$ scale nearly linearly with $\lambda$. Moreover, the inset of Fig. 3(d) shows that in the Dicke limit the EE grows linearly at early times, supporting its role as a sensitive probe of transient chaos.
@@
The total Hamiltonian is given by ^=^$\hat{\mathcal{H}}=\hat{\mathcal{H}}_{\rm TD}+\frac{\mu}{\sqrt{N}}\hat{\mathcal{H}}_{\rm I}$, where ^$\hat{\mathcal{H}}_{\rm TD}$ is a tridiagonal $N\times N$ random matrix and ^$\hat{\mathcal{H}}_{\rm I}$ is a perturbation whose structure we vary to explore different dynamical regimes. The parameter $\mu\in[0,1]$ controls the relative strength of the perturbation. Dissipation is incorporated through a random jump operator $\hat{L}$, which has nonzero real random entries only on the first subdiagonal, $\hat{L}_{m+1,m}\neq 0$. The resulting Liouvillian superoperator takes the standard Lindblad form,
$\hat{\mathcal{L}}=-$ $\,i\left[\hat{\mathcal{H}}\otimes\mathbb{I}-\mathbb{I}\otimes\hat{\mathcal{H}}\right]$
+\,+ $\gamma\Big{(}2\hat{L}\otimes\hat{L}^{*}-\hat{L}^{\dagger}\hat{L}\otimes\mathbb{I}-\mathbb{I}\otimes\hat{L}^{\mathrm{T}}\hat{L}^{*}\Big{)},$ (2)
where $\gamma$ is the dissipation strength.
@@
We first choose ^$\hat{\mathcal{H}}_{\rm I}=\hat{\mathcal{H}}_{\rm GOE}$, that is, a full random matrix drawn from the Gaussian orthogonal ensemble. In the unitary case ($\gamma=0$), the spectral statistics of $\hat{\mathcal{H}}$ interpolates from Poisson ($\mu=0$) to Wigner-Dyson ($\mu=1$) level spacing distribution. When dissipation is included ($\gamma=1$), the Liouvillian spectrum transitions from the two-dimensional (2d) Poisson ($\mu=0$) to the Ginibre ($\mu=1$) statistics of non-Hermitian random matrices [10, 19, 20, 18], as shown in the SM [54].

### v2

@@ Abstract
Dissipative quantum chaos plays a central role in the characterization and control of information scrambling, non-unitary evolution, and thermalization, but it still lacks a precise definition. The Grobe-Haake-Sommers conjecture, which links Ginibre level repulsion to classical chaotic dynamics, was recently shown to fail [Phys. Rev. Lett. 133, 240404 (2024)]. We properly restore the quantum-classical correspondence through a dynamical approach based on the von Neumann entropy (VNE) and out-of-time-order correlators (OTOCs), which reveal signatures of chaos beyond spectral statistics. Focusing on the open anisotropic Dicke model, we identify two distinct regimes: transient chaos, marked by rapid early-time growth of VNE and OTOCs followed by low saturation values, and steady-state chaos, characterized by high long-time values. We introduce a random matrix toy model and show that Ginibre spectral statistics signals short-time chaos rather than steady-state chaos. Our results establish VNE dynamics and OTOCs as reliable diagnostics of dissipative quantum chaos across different timescales.
@@
Quantum chaos is connected to a wide range of theoretical and experimental phenomena, from thermalization [1, 2] and the failure of many-body localization [3, 4], to the scrambling of quantum information [5, 6] and the exponential growth of perturbations near black hole horizons [7]. While chaos in closed quantum systems is typically characterized by random matrix theory [8, 9], including Wigner-Dyson level statistics [10, 11, 12] and eigenstates resembling random vectors [13, 1], defining chaos in open quantum systems remains a challenge [14, 15, 16]. Dissipation alters dynamical behavior, raising questions about the quantum-classical correspondence in non-unitary settings.
@@
Efforts to define quantum chaos in open systems have led to the Grobe-Haake-Sommer conjecture [17, 18, 11, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], which relates classical chaotic attractors with Ginibre level repulsion in the spectrum of the Liouvillian superoperator. However, recent studies have shown that this correspondence can fail [29, 30, 31, 32]. In particular, Ginibre spectral statistics may emerge even when the long-time dynamics is regular [29]. This discrepancy suggests that spectral indicators alone may be insufficient to capture the full nature of dissipative quantum chaos and dynamical properties should also be considered [30].
@@
Figure 1: Schematic illustration. A dynamical framework restores the quantum-classical correspondence by identifying three regimes: (I) steady-state chaos, featuring a chaotic attractor in the classical dynamics (left), and linear growth followed by large saturation of the von Neumann entropy (middle panel); (II) transient chaos, characterized by a regular attractor, and rapid short-time VNE growth followed by decay at long times; and (III) regular dynamics, showing slow VNE growth and low saturation. Regimes (I) and (II) exhibit Ginibre Liouvillian spectral statistics (right panels), while (III) corresponds to 2d-Poisson statistics.
In this work, we adopt a dynamical perspective and demonstrate that the quantum-classical correspondence can be restored when chaos is diagnosed through the evolution of the von Neumann entropy (VNE) and out-of-time-order correlators (OTOCs). We generalize fidelity OTOCs (FOTOCs) for non-unitary evolution, making them applicable to open quantum systems. Unlike closed Hamiltonian systems, where rapid early-time mixing often indicates persistent chaotic dynamics, dissipation can suppress long-time chaos, despite signs of chaotic behavior at short times. As illustrated schematically in Fig. 1, we identify three regimes with qualitatively distinct behavior: (I) steady-state chaos, marked by fast initial growth and large long-time saturation values of VNE (middle panel), in agreement with the presence of a chaotic attractor in the classical dynamics (see Bloch sphere on the left side); (II) transient chaos, which also features rapid early-time scrambling but evolves into a low-entropy steady state, consistent with the presence of a stable attractor; and (III) regular regime, with slow dynamics and no chaos at any time. The FOTOCs exhibit the same qualitative dynamical behavior across the three regimes, showing rapid early-time growth in both cases (I) and (II). Importantly, the right panels of Fig. 1 show that both transient and steady-state chaos exhibit Ginibre level statistics in the Liouvillian spectrum, demonstrating the limitations of spectral features in fully characterizing dissipative quantum chaos.
@@
We explore these ideas using the anisotropic Dicke model (ADM) [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54] in the presence of photon loss. We show that, unlike Liouvillian spectral statistics [29], the dynamics of the VNE of a subsystem of the Dicke model and FOTOCs distinguish transient from steady-state chaos, in agreement with classical diagnostics based on Lyapunov exponents.
@@
To support and generalize our findings, we introduce a random matrix toy model with a tunable Liouvillian. We show that Ginibre spectral statistics emerge whenever the VNE exhibits rapid initial linear growth, regardless of whether it saturates at high values (steady-state chaos) or decays to low values (transient chaos). This analysis reveals that Liouvillian spectral statistics can reflect early-time chaotic behavior, but does not distinguish different asymptotic regimes.
@@
[…]
where $\hbar=1$, $\hat{a}$ ($\hat{a}^{\dagger}$) annihilates (creates) a cavity photon mode with frequency $\omega$, the collective pseudospin operators $\hat{S}_{z,+,-}$ describe the joint behavior of $N$ two-level atoms with energy splitting $\omega_{0}$, and $\lambda_{\pm}$ are the atom-photon coupling strengths. In all figures, we set $\omega=1.0$, $\omega_{0}=1.0$, and $S=N/2=5$ (results for larger $S$ are in the Supplemental Material (SM) [55]).
@@
[…]
@@
[…]
@@
Steady-state chaos.– In the open classical ADM, chaotic dynamics emerges when the regular attractors become unstable. In the quantum domain, the steady state is defined by the zero eigenvalue of the Liouvillian superoperator  [^1] Despite the presence of a limit cycle in the open classical systems, its quantum analogue can only be probed in the asymptotic limit $S\to\infty$, where some eigenvalues of the Liouvillian become purely imaginary [84, 85, 86].. We show that this long-time quantum state can retain signatures of the underlying classical chaos, a phenomenon we refer to as steady-state chaos.
@@
We investigate the time evolution of the total VNE defined as $\mathcal{S}^{\rm VN}(t)=\mathcal{S}_{\rm spin}^{\rm VN}(t)+\mathcal{S}_{\rm photon}^{\rm VN}(t)$, where $\mathcal{S}_{\rm spin}^{\rm VN}$ ($\mathcal{S}_{\rm photon}^{\rm VN}$) is the VNE of the reduced density matrix of the spin (photon) subsystem obtained by tracing out the photon (spin) degrees of freedom. In the absence of dissipation, these entropies reduce to the entanglement entropy, which builds up dynamically due to the atom–photon interaction and is commonly used to diagnose chaos in isolated systems [2, 1, 60, 61, 62, 63]. The open system is initially prepared in a product state consisting of a spin coherent state and a photon coherent state [64].
@@
Contrary to Hamiltonian systems, where the long-time saturation value of the VNE of subsystems depends on the initial state, dissipation drives the system toward a unique asymptotic behavior. In dissipative dynamics, only the short-time evolution of the VNE depends on the initial state.
@@
Figure 2: Signature of steady-state chaos on the $\lambda_{-}-\lambda_{+}$ plane. (a), (c) Averaged long-time Lyapunov exponent $\Lambda_{\rm ss}$ and (b), (d) total von Neumann entropy of the spin and photon subsystems $\mathcal{S}_{\rm ss}^{\rm VN}$ in the (a)-(b) absence of dissipation, $\kappa=0$, and (c)-(d) presence of dissipation, $\kappa=1$.
@@
The color plots in Fig. 2 compare the average long-time Lyapunov exponent [65, 66, 67], denoted by $\Lambda_{\rm ss}$ [Figs. 2(a),2(c)], with the steady-state total VNE, $\mathcal{S}_{\rm ss}^{\rm VN}$, averaged over an ensemble of initial states [Figs. 2(b),2(d)] for the isolated [Figs. 2(a)-(b)] and dissipative [Figs. 2(c)-(d)] cases. In the absence of dissipation, the system exhibits large values of $\Lambda_{\rm ss}$ and $\mathcal{S}_{\rm ss}^{\rm VN}$ over a broad region of the coupling parameter space, especially near the Dicke limit, $\lambda_{-}=\lambda_{+}=\lambda$, where the classical dynamics is maximally chaotic. When dissipation is introduced, the extent of the chaotic region is significantly reduced. As seen in Figs. 2(c)-(d), large values of $\mathcal{S}_{\rm ss}^{\rm VN}$ persist only within a narrow triangular region of the parameter space, where chaos is identified classically by positive values of $\Lambda_{\rm ss}$. Outside this region, the classical system becomes regular [see Fig. 2(c)], including the area around the Dicke limit, and in the quantum case, small but nonzero values of $\mathcal{S}_{\rm ss}^{\rm VN}$ remain visible [Fig. 2(d)], due to quantum fluctuations at finite spin $S$ and the proximity to a non-equilibrium phase transition. Although the magnitude of $\mathcal{S}_{\rm ss}^{\rm VN}$ does not quantify the absolute degree of chaos, as the Lyapunov exponent does, it efficiently distinguishes between chaotic and regular regimes, with this distinction becoming more pronounced as $S$ increases (see [55]).
@@
The key result in Fig. 2 is the strong quantum-classical correspondence observed not only in the isolated regime but also in the presence of dissipation. The saturation value of the VNE closely follows the classical Lyapunov exponent, as shown by the close agreement between Figs. 2(c) and 2(d). This demonstrates that, unlike Liouvillian level statistics, which can present Ginibre spectral correlations when the open classical model is regular [29], the steady-state VNE provides a reliable quantum signature of dissipative classical chaos.
@@
In isolated systems, the OTOC has also been extensively studied as a quantum diagnostic of chaos [68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 43, 44, 81, 82, 83]. It is defined as $\mathcal{F}(t)=\langle\hat{W}^{\dagger}(t)\hat{V}^{\dagger}(0)\hat{W}(t)\hat{V}(0)\rangle$, where the operator $\hat{W}$ evolves in time and the operator $\hat{V}$ is fixed at $t=0$. The OTOC quantifies the spread of quantum information and sensitivity to perturbations. A variant of this correlator is the FOTOC [80] obtained by choosing $\hat{V}=\hat{\rho}(0)$ and $\hat{W}=e^{i\delta\phi\hat{G}}$, where $\delta\phi$ is a small perturbation and $\hat{G}$ is a Hermitian operator. In the perturbative limit $\delta\phi\ll 1$, the FOTOC can be written in terms of the variance $(\Delta{G}(t))^{2}=\langle\hat{G}^{2}(t)\rangle-\langle\hat{G}(t)\rangle^{2}$ as $\mathcal{F}_{G}(t)\approx 1-\delta\phi^{2}(\Delta{G}(t))^{2}$, and its early-time exponential growth indicates instability and chaos.
@@
We generalize the FOTOC to open quantum systems by evolving the operator $\hat{W}$ under non-unitary dynamics according to $d\hat{W}/dt=\mathcal{L}^{\dagger}[\hat{W}$]. Similar to the VNE, the steady-state value of the standard deviation $(\Delta G)_{\rm ss}$ obtained from the FOTOC also serves as an indicator of steady-state chaos (see SM [55]).
@@
Transient chaos.– Complementing the long-time analysis, we examine the short-time dynamics of the VNE and $\Delta G(t)$, both of which capture signatures of transient chaos. In isolated classical systems, chaos arises from rapid mixing and sensitivity to initial conditions that emerge early in the evolution. When dissipation is introduced, the system may still display chaotic behavior at short times, even if the long-time dynamics become regular. This transient chaos reflects a competition between the early-time chaotic dynamics inherited from the isolated system and the eventual convergence to a regular attractor of the dissipative system. In the quantum regime, a similar phenomenon occurs: the dynamics can ultimately settle into a non-chaotic steady state, even if remnants of chaos, such as rapid entropy growth, are still visible at early times. In our system, the Dicke limit offers a natural setting to explore this behavior, as it exhibits maximal chaos in the absence of dissipation, while increasing the dissipation rate $\kappa$ progressively suppresses chaos at long times, as the superradiant fixed point becomes a stable attractor.
@@
Figure 3: Transient chaos vs steady-state chaos in the Dicke limit $\lambda_{-}=\lambda_{+}=\lambda$. (a) Finite-time ensemble averaged Lyapunov exponent $\Lambda_{t}$, (b) total VNE $\mathcal{S}^{\rm VN}(t)$, and (c) standard deviation $\Delta S_{z}(t)\approx\sqrt{1-\mathcal{F}_{z}(t)}/\delta\phi$ obtained from FOTOC $\mathcal{F}_{z}(t)$ of $z$ component of spin for $\delta\phi\ll 1/S$, for atom-photon coupling strengths $\lambda=1.2,2.0$ and $\kappa=0,1$. In (b), the time axis is linear for $t\leq 0.5$ and logarithmic for $t>0.5$. (d) The linear growth rate of VNE, $\mathcal{S}^{\rm VN}_{\rm slope}$, and the time-averaged Lyapunov exponent, $\overline{\Lambda}$, within a time interval $t\in[0,0.5]$ as a function of the coupling strength $\lambda$ for $\kappa=1$.
@@
To uncover quantum signatures of transient chaos in the Dicke limit, we examine the time evolution of the von Neumann entropy $\mathcal{S}^{\rm VN}(t)$ [Fig. 3(b)] and the standard deviation of the $z$ component of the spin operator, $\Delta S_{z}(t)$ [Fig. 3(c)], both averaged over an ensemble of initial states. In both isolated and dissipative settings, these observables exhibit rapid growth at short times, consistent with short-time chaotic dynamics. However, at long times, similar to the behavior of $\Lambda_{t}$, the asymptotic values of $\mathcal{S}^{\rm VN}(t)$ and $\Delta S_{z}(t)$ are significantly reduced for the dissipative model, indicating a transition to regular dynamics. This contrast between rapid short-time growth and long-time suppression reflects the presence of transient chaos also for the quantum dissipative system. Similar behavior is also observed for the standard deviation of the photon number.
Further insight into the nature of transient chaos is provided by its dependence on the atom-photon coupling $\lambda$. As seen in Figs. 3(b)-(c), the short-time growth rates of $\mathcal{S}^{\rm VN}(t)$ and $\Delta S_{z}(t)$ increase with the atom-photon coupling $\lambda$, in agreement with the larger values of the short-time Lyapunov exponent observed in Fig. 3(a). This dependence on $\lambda$ is quantified in Fig. 3(d), where we show that both the growth rate of VNE, $\mathcal{S}^{\rm VN}_{\rm slope}$, and the time-averaged Lyapunov exponent computed over small time windows $t\in[0,0.5]$ scale nearly linearly with $\lambda$. Moreover, Fig. 3(b) shows that in the Dicke limit, the VNE grows linearly at early times, supporting its role as a sensitive probe of transient chaos.
@@
The total Hamiltonian is given by $\hat{\mathcal{H}}=\hat{\mathcal{H}}_{\rm TD}+\frac{\mu}{\sqrt{N}}\hat{\mathcal{H}}_{\rm I}$, where $\hat{\mathcal{H}}_{\rm TD}$ is a tridiagonal $N\times N$ random matrix and $\hat{\mathcal{H}}_{\rm I}$ is a perturbation whose structure we vary to explore different dynamical regimes. The parameter $\mu\in[0,1]$ controls the relative strength of the perturbation. Dissipation is incorporated through a random jump operator $\hat{L}$, which has nonzero real random entries only on the first subdiagonal, $\hat{L}_{m+1,m}\neq 0$. The resulting Liouvillian superoperator takes the standard Lindblad form,
$\hat{\mathcal{L}}$=- $\,i\left[\hat{\mathcal{H}}\otimes\mathbb{I}-\mathbb{I}\otimes\hat{\mathcal{H}}^{*}\right]$
$+\,$ $\gamma\Big(2\hat{L}\otimes\hat{L}^{*}-\hat{L}^{\dagger}\hat{L}\otimes\mathbb{I}-\mathbb{I}\otimes\hat{L}^{\mathrm{T}}\hat{L}^{*}\Big),$ (2)
where $\gamma$ is the dissipation strength and “T” represents the transposition of a matrix.
@@
We first choose $\hat{\mathcal{H}}_{\rm I}=\hat{\mathcal{H}}_{\rm GOE}$, that is, a full random matrix drawn from the Gaussian orthogonal ensemble. In the unitary case ($\gamma=0$), the spectral statistics of $\hat{\mathcal{H}}$ interpolates from Poisson ($\mu=0$) to Wigner-Dyson ($\mu=1$) level spacing distribution. When dissipation is included ($\gamma=1$), the Liouvillian spectrum transitions from the two-dimensional (2d) Poisson ($\mu=0$) to the Ginibre ($\mu=1$) statistics of non-Hermitian random matrices [11, 21, 22, 20], as shown in the SM [55].
To probe the resulting dynamics of the open system ($\gamma=1$), we initialize it in a pure product state, $\hat{\rho}(0)$, built with the eigenstates of $\hat{\mathcal{H}}_{\rm TD}$, and evolve it under the Liouvillian in Eq. (2). We assume that the dimension $N$ of the total Hamiltonian is constructed as the tensor product of two abstract subsystems, each of dimension $M$, so $N=M^{2}$.